DIRC Operations Weekly Report

Period: Saturday 2006/06/03 - Friday 2006/06/09

Oncall: Emmanuel Latour, Rolf Andreassen
Fast Monitoring: Emmanuel Latour, Rolf Andreassen

File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20060609.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20060609 (empty)

Contents


BaBar-PEP news

PEP peak lumi: 10.3*1033 (previous week: 10.4 1033)

BaBar integrated lumi: 2.943 fb-1 (previous week: 2.669 fb-1)

Run 5b PEP delivery: 55.53 fb-1

Run 5 total delivery: 113.97 fb-1 Run 5 total BaBar recorded: 109.72 fb-1 including 6.26 fb-1 off-peak data

PEP activities this week:

  • Tue morning: characterisation of the machine
  • Wed: PEP PR02 micro problem+LER region 4 vacuum problem: had to remove a chamber and braise it. Beams back on Saturday (June, 10th) owl shift.
  • Thu: a klystron failed during swing shift due to a water leak. They replaced it.

    PEP activities planned next week :

  • Carry on offpeak data taking

    Foreseen work opportunities (PEP down periods): None.

    BaBar activities/issues this week: UPS system tripped on Saturday owl shift (at about 4:30pm). Lasted 11 hours.

  • Loss of power in almost all EH.
  • Loss of monitoring/computing
  • Pump for cooling water turned off
  • Failuer of SVT chillers
  • UPS system finally bypassed (brought back within BaBar on Thursday while magnet was ramped down)

    DIRC activities

    See DIRC logbook for details

    Downtime due to the DIRC: None.

    Backgrounds

    Background summary plot

    Scheduled activities

    Pending Work

    Links